ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

9. Ordinary Differential Equations and Dynamical Systems

1991 MS Classification: 34 Yebdri, Mustapha, Université de Tlemecen, Algeria: The linear stability of slowly oscillating periodic solutions of a delay differential equation. The equation x′(t) = −μx(t) + f(x(t − 1)), with μ0 and a smooth function f : R → R satisfying f(0) = 0 and xf((x) < 0∀x ≠ 0, has applications in many domains. Slowly oscillating solutions are very important in the study of the dynamic of this equation. A result on the linear stability of such periodic solutions will be presented. Reference: M. Yebdri, Ph. D Ghesis, University of Giessen, Germany(1996).